You currently have $3,900 (Present Value) in an account that has an interest rate of 4.5% per year compounded semi-annually (2 times per year). You want to withdraw all your money when it reaches $9,360 (Future Value). In how many years will you be able to withdraw all your money?
Answer #1

Semi annual rate = 4.5% / 2 = 2.25%

Number of periods = 2*n

Future value = Presnet value (1 + r)n

9,360 = 3,900 (1 + 0.0225)2*n

2.4 = (1.0225)2n

LN 2.4 = 2n LN 1.0225

0.875469 = 2n*0.022251

0.875469 = n*0.044501

n = 19.67

It will take 19.67 years